# Transverse Instability of Stokes Waves at Finite Depth

### v1

@@ 4.2. Third-order expansions of $\textrm{L}_{\varepsilon,\delta}$
$U^{\varepsilon,\delta}_{j}$ =$=\{1-(P_{\varepsilon,\delta}-P_{0,0})^{2}\}^{-\frac{1}{2}}P_{\varepsilon,\delta}U_{j}.$
@@
A similar argument shows that $q(\xi)>0$, and in [17] we show that $r(\xi)>0$ and $21<\xi<$ 2. It follows that
@@
Given the behavior of $b_{3,0}$ as $h\rightarrow 0^{+}$ and $h\rightarrow\infty$, $b_{3,0}$ must have at least one zero in $(0,\infty)$. Since $b_{3,0}$ is analytic on $(0,\infty)$, its zero set $\mathcal{Z}$ cannot have an accumulation point in $(0,\infty)$. $\mathcal{Z}$ is bounded because $\lim_{h\to\infty}b_{3,0}(h)<0$. Moreover, 00 cannot be an accumulation point of $\mathcal{Z}$ because $\lim_{h\to 0^{+}}b_{3,0}(h)=\infty$. Thus $\mathcal{Z}$ is finite.
[…]
@@
where $A+C$ is the trace of $\text{L}_{\varepsilon,\delta}-i\sigma I$ and so is real-valued. Since $A$, $B$, and $C$ are real analytic in $(\varepsilon,\delta)$ in a neighborhood of $(0,0)$, so are $A+C$ and $\Delta(\varepsilon,\delta)$. Moreover, it follows from the expansions of $A$, $B$, and $C$ that $\Delta(\varepsilon,\delta)=O(\delta^{2})$ and $A+C=O(\delta)$ as $(\varepsilon,\delta)\rightarrow(0,0)$. We will complete the proof of Theorem 1.1 by showing that the characteristic polynomial (5.5) has a root with positive real part, that is, $\Delta(\varepsilon,\delta)>0$ for suitably small $\varepsilon$ and $\delta$.
@@
$\delta=\kappa_{0}\varepsilon^{2}+\theta\varepsilon^{3},\quad\textrm{where}\quad\kappa_{0}:=\frac{a_{2,0}-c_{2,0}}{a_{0,1}-c_{0,1}}.$ (5.8)
@@
$|\theta|<\kappa_{1},\quad\textrm{where}\quad\kappa_{1}:=\frac{2|b_{3,0}|}{|a_{0,1}-c_{0,1}|},$ (5.10)
@@
±$\qquad\pm\frac{1}{2}\left[4b_{3,0}^{2}-(a_{0,1}-c_{0,1})^{2}\theta^{2}\right]^{\frac{1}{2}}|\varepsilon|^{3}+O(\varepsilon^{4}).$
@@
Appendix A Normal trace of vector fields
We recall the following classical trace operator.
Proposition A.1 (Section 3.2, Chapter IV, [9]).
$H_{\operatorname{div}}(U)=\{u\in L^{2}(U;\mathbb{R}^{n}):\operatorname{div}u\in L^{2}(U)\}.$ (A.1)
There exists a bounded linear operator $\gamma:H_{\operatorname{div}}(U)\to H^{\frac{1}{2}}(\partial U)$ such that $\gamma(u)=u\cdot\nu$ for $u\in C^{\infty}_{0}(\overline{U})$, where $\nu$ is the unit outward normal to $\partial U$. In particular, there exists a constant $C>0$ such that
$\|\gamma(u)\|_{H^{-\frac{1}{2}}(\partial U)}\leq C\Big{(}\|u\|_{L^{2}(U)}+\|\operatorname{div}u\|_{L^{2}(U)}\Big{)}\quad\forall u\in H_{\operatorname{div}}(U).$ (A.2)
$\langle\gamma(u),\gamma_{0}(w)\rangle_{H^{-\frac{1}{2}}(\partial U),H^{\frac{1}{2}}(\partial U)}=\int_{U}u\cdot\nabla w+\int_{U}w\operatorname{div}u$ (A.3)
holds, where $\gamma_{0}(w)$ is the trace of $w$.
The following lemma provides an extension of the preceding trace operator for a cylindrical shell that is sufficient for our purposes.
Lemma A.2.
Let $U=\mathbb{T}\times(a,b)$, where $-\infty<a<b<\infty$. Since $\partial U=(\mathbb{T}\times\{a\})\cup(\mathbb{T}\times\{b\})$, we identify any function $g$ on $\partial U$ with the pair $(g_{1},g_{2})\equiv(g(\cdot,a),g(\cdot,b)):\mathbb{T}\to\mathbb{R}^{2}$. We denote a point in $U$ by $(x,z)$ and set
$\tilde{H}_{\operatorname{div}}(U)=\{u\in L^{2}(U;\mathbb{R}^{2}):\operatorname{div}u=\partial_{x}^{2}F\quad\text{for~{}some~{}}F\in H^{1}(U)\}.$ (A.4)
(i) There exists a linear operator $\tilde{\gamma}:\tilde{H}_{\operatorname{div}}(U)\to(H^{-\frac{1}{2}}(\mathbb{T}))^{2}$ with the following properties
  • (a)
    $\tilde{\gamma}(u)=\gamma(u)$ if $u\in\tilde{H}_{\operatorname{div}}(U)\cap H_{\operatorname{div}}(U)$.
  • (b)
    For some constant $C=C(a,b)>0$, we have
    $\|\tilde{\gamma}(u)\|_{(H^{-\frac{1}{2}}(\mathbb{T}))^{2}}\leq C(\|u\|_{L^{2}(U)}+\|\partial_{x}F\|_{L^{2}(U)}),$ (A.5)
    where the right-hand side of (A.5) is independent of the choice of $F$.
  • (c)
    If $u=\begin{bmatrix}u_{1}\\ 0\end{bmatrix}\in\tilde{H}_{\operatorname{div}}(U)$, then $\tilde{\gamma}(u)=0$.
(ii) We extend $\tilde{\gamma}$ to $\tilde{H}_{\operatorname{div}}(U)+H_{\operatorname{div}}(U)$ by defining
$\tilde{\gamma}(u)=\tilde{\gamma}(\tilde{v})+\gamma(v)$ (A.6)
if $u=\tilde{v}+v$, $\tilde{v}\in\tilde{H}_{\operatorname{div}}(U)$, $v\in H_{\operatorname{div}}(U)$. This definition is independent of the choice of the pair $(\tilde{v},v)$. As a consequence, if $u\in L^{2}(U;\mathbb{R}^{2})$ and $\operatorname{div}u=G+\partial_{x}^{2}F$ for some $G\in L^{2}(U)$ and $F\in H^{1}(U)$, then
$\|\tilde{\gamma}(u)\|_{(H^{-\frac{1}{2}}(\mathbb{T}))^{2}}\leq C(\|u\|_{L^{2}(U)}+\|G\|_{L^{2}(U)}+\|\partial_{x}F\|_{L^{2}(U)}).$ (A.7)
Proof.
(i) For any $u\in\tilde{H}_{\operatorname{div}}(U)$ with $\operatorname{div}u=\partial_{x}^{2}F$, using a lifting operator $\ell_{U}:(H^{\frac{1}{2}}(\mathbb{T}))^{2}\to H^{1}(U)$, we define $\tilde{\gamma}(u)\in(H^{-\frac{1}{2}}(\mathbb{T}))^{2}$ by
$\langle\tilde{\gamma}(u),g\rangle_{(H^{-\frac{1}{2}}(\mathbb{T}))^{2},(H^{\frac{1}{2}}(\mathbb{T}))^{2}}:=(u,\nabla\ell_{U}g)_{L^{2}(U)}-(\partial_{x}F,\partial_{x}\ell_{U}g)_{L^{2}(U)}\quad\forall g\in(H^{\frac{1}{2}}(\mathbb{T}))^{2}.$ (A.8)
Clearly (A.8) defines $\tilde{\gamma}(u)$ as a bounded linear functional on $(H^{\frac{1}{2}}(\mathbb{T}))^{2}$ with norm bounded by $C(\|u\|_{L^{2}(U)}+\|\partial_{x}F\|_{L^{2}(U)})$. Thus (A.5) is valid. This norm bound is independent of the choice of $F$. Indeed, if $\operatorname{div}u=\partial_{x}^{2}F=\partial_{x}^{2}F_{1}$, then $F(x,z)-F_{1}(x,z)$ is independent of $x$ due to the periodicity in $x$, whence $\partial_{x}F=\partial_{x}F_{1}$.
Now suppose that $u\in\tilde{H}_{\operatorname{div}}(U)\cap H_{\operatorname{div}}(U)$, so that $L^{2}(U)\ni\operatorname{div}u=\partial_{x}^{2}F$ for some $F\in H^{1}(U)$. The Stokes formula (A.3) implies
$\langle\gamma(u),g\rangle_{H^{-\frac{1}{2}}(\partial U),H^{\frac{1}{2}}(\partial U)}$ $=(u,\nabla\ell_{U}g)_{L^{2}(U)}+(\partial^{2}_{x}F,\ell_{U}g)_{L^{2}(U)}.$
Since $\partial_{x}^{2}F=\operatorname{div}u\in L^{2}(U)$, $F\in H^{1}(U)$, and $\ell_{U}g\in H^{1}(U)$, we have $(\partial^{2}_{x}F,\ell_{U}g)_{L^{2}(U)}=-(\partial_{x}F,\partial_{x}\ell_{U}g)_{L^{2}(U)}$ by the periodicity in $x$, and hence for all $g\in H^{\frac{1}{2}}(\mathbb{T}))^{2}$,
$\langle\gamma(u),g\rangle_{H^{-\frac{1}{2}}(\partial U),H^{\frac{1}{2}}(\partial U)}=(u,\nabla\ell_{U}g)_{L^{2}(U)}-(\partial_{x}F,\partial_{x}\ell_{U}g)_{L^{2}(U)}=\langle\tilde{\gamma}(u),g\rangle_{(H^{-\frac{1}{2}}(\mathbb{T}))^{2},(H^{\frac{1}{2}}(\mathbb{T}))^{2}}.$
Therefore, $\tilde{\gamma}=\gamma$ on $\tilde{H}_{\operatorname{div}}(U)\cap H_{\operatorname{div}}(U)$.
Assume that $u=\begin{bmatrix}u_{1}\\ 0\end{bmatrix}\in\tilde{H}_{\operatorname{div}}(U)$. Then $\partial_{x}^{2}F=\operatorname{div}u=\partial_{x}u_{1}$ for some $F\in H^{1}(U)$, and hence $u_{1}(x,z)=\partial_{x}F(x,z)+f(z)$ for some $f\in L^{2}((a,b))$. The definition (A.8) implies that for all $g\in(H^{\frac{1}{2}}(\mathbb{T}))^{2}$ we have
$\langle\tilde{\gamma}(u),g\rangle_{(H^{-\frac{1}{2}}(\mathbb{T}))^{2},(H^{\frac{1}{2}}(\mathbb{T}))^{2}}$ $=(u_{1},\partial_{x}\ell_{U}g)_{L^{2}(U)}-(\partial_{x}F,\partial_{x}\ell_{U}g)_{L^{2}(U)}$
$=(f,\partial_{x}\ell_{U}g)_{L^{2}(U)}=\int_{a}^{b}f(z)\int_{\mathbb{T}}\partial_{x}\ell_{U}g(x,z)dxdz=0$
due to the periodicity in $x$. Thus $\tilde{\gamma}(u)=0$.
(ii) Next, we show that the definition (A.6) of the extended $\tilde{\gamma}$ is well defined. Suppose that $u=\tilde{v}+v=\tilde{w}+w$, where $\tilde{v}$, $\tilde{w}\in\tilde{H}_{\operatorname{div}}(U)$, and $v$, $w\in H_{\operatorname{div}}(U)$. Then $\tilde{v}-\tilde{w}=w-v\in\tilde{H}_{\operatorname{div}}(U)\cap H_{\operatorname{div}}(U)$, and thus $\tilde{\gamma}(\tilde{v}-\tilde{w})=\gamma(w-v)$. It follows that $\tilde{\gamma}(\tilde{v})+\gamma(v)=\tilde{\gamma}(\tilde{w})+\gamma(w)$.
Suppose that $u\in L^{2}(U;\mathbb{R}^{2})$ and $\operatorname{div}u=G+\partial_{x}^{2}F$ for some $G\in L^{2}(U)$ and $F\in H^{1}(U)$. Let $\phi:U\to\mathbb{R}$ be the $H^{2}(U)\cap H^{1}_{0}(U)$ solution to the Poisson equation $\Delta\phi=G$, and set $v=\nabla\phi$, $\tilde{v}=u-v$. Then $v\in H_{\operatorname{div}}(U)$, $\tilde{v}\in\tilde{H}_{\operatorname{div}}(U)$ with $\operatorname{div}\tilde{v}=\partial_{x}^{2}F$, and $\|\phi\|_{H^{2}(U)}\leq C\|G\|_{L^{2}(U)}$. Therefore it follows from (A.6), (A.5) and (A.2) that
$\|\tilde{\gamma}(u)\|_{(H^{-\frac{1}{2}}(\mathbb{T}))^{2}}$ $\leq\|\tilde{\gamma}(\tilde{v})\|_{(H^{-\frac{1}{2}}(\mathbb{T}))^{2}}+\|\gamma(v)\|_{(H^{-\frac{1}{2}}(\mathbb{T}))^{2}}$
$\leq C\left(\|\tilde{v}\|_{L^{2}(U)}+\|\partial_{x}F\|_{L^{2}(U)}+\|v\|_{L^{2}(U)}+\|\operatorname{div}v\|_{L^{2}(U)}\right)$
$\leq C\left(\|u\|_{L^{2}(U)}+\|v\|_{L^{2}(U)}+\|\partial_{x}F\|_{L^{2}(U)}+\|v\|_{L^{2}(U)}+\|\Delta\phi\|_{L^{2}(U)}\right),$
which implies (A.7). ∎

### v2

@@ 4.2. Third-order expansions of $\textrm{L}_{\varepsilon,\delta}$
$U^{\varepsilon,\delta}_{j}$ $=\{1-(P_{\varepsilon,\delta}-P_{0,0})^{2}\}^{-\frac{1}{2}}P_{\varepsilon,\delta}U_{j}.$
@@
A similar argument shows that $q(\xi)>0$, and in [18] we show that $r(\xi)>0$ and $1<\xi<2$. It follows that
@@
Given the behavior of $b_{3,0}$ as $h\rightarrow 0^{+}$ and $h\rightarrow\infty$, $b_{3,0}$ must have at least one zero in $(0,\infty)$. Since $b_{3,0}$ is analytic on $(0,\infty)$, its zero set $\mathcal{Z}$ cannot have an accumulation point in $(0,\infty)$. $\mathcal{Z}$ is bounded because $\lim_{h\to\infty}b_{3,0}(h)<0$. Moreover, 0 cannot be an accumulation point of $\mathcal{Z}$ because $\lim_{h\to 0^{+}}b_{3,0}(h)=\infty$. Thus $\mathcal{Z}$ is finite.
Although by analytic reasoning we have not been able to determine the number of zeros of $b_{3,0}$, a numerical calculation indicates that there is exactly one. Indeed, the numerical plot of $b_{3,0}$ in Figure 3 reveals that exactly one such depth occurs at $h_{crit}=0.25065$.... In order to obtain Figure 3, we use Newton’s method to compute $\beta_{*}(h)$ for a given $h$ from the resonance condition (3.4). This in turn allows us to evaluate $b_{3,0}$ numerically at depth $h$. Repeating this procedure for $0.1<h<4$ in step sizes of $\Delta h=0.01$ yields Figure 3. The critical depth $h_{crit}=0.25065$... is obtained by bisection using increasing smaller $\Delta h$ near the critical depth.
@@
where we recall that $\sigma+\frac{1}{2}(A+C)$ is real. Since $A$, $B$, and $C$ are real analytic in $(\varepsilon,\delta)$ in a neighborhood of $(0,0)$, so are $A+C$ and $\Delta(\varepsilon,\delta)$. Moreover, it follows from the expansions of $A$, $B$, and $C$ that $\Delta(\varepsilon,\delta)=O(\delta^{2})$ and $A+C=O(\delta)$ as $(\varepsilon,\delta)\rightarrow(0,0)$. We will complete the proof of Theorem 1.1 by showing that the characteristic polynomial (5.5) has a root with positive real part, that is, $\Delta(\varepsilon,\delta)>0$ for suitably small $\varepsilon$ and $\delta$.
@@
$\delta=\kappa_{0}\varepsilon^{2}+\theta\varepsilon^{3},\qquad\textrm{where}\qquad\kappa_{0}:=\frac{a_{2,0}-c_{2,0}}{a_{0,1}-c_{0,1}}.$ (5.8)
@@
$|\theta|<\kappa_{1},\qquad\textrm{where}\qquad\kappa_{1}:=\frac{2|b_{3,0}|}{|a_{0,1}-c_{0,1}|},$ (5.10)
@@
$\qquad\pm\frac{1}{2}\left[4b_{3,0}^{2}-(a_{0,1}-c_{0,1})^{2}\theta^{2}\right]^{\frac{1}{2}}|\varepsilon|^{3}+O(\varepsilon^{4}).$